### v1

@@ Abstract
Self-supervised pretraining from static structures of drug-like compounds and proteins enable powerful learned feature representations. Learned features demonstrate state of the art performance on a range of predictive tasks including molecular properties, structure generation, and protein-ligand interactions. The majority of approaches are limited by their use of static structures and it remains an open question, how best to use atomistic molecular dynamics (MD) simulations to develop more generalized models to improve prediction accuracy for novel molecular structures. We present SURrogate mmGBSA (SurGBSA) as a new modeling approach for MD-based representation learning, which learns a surrogate function of the Molecular Mechanics Generalized Born Surface Area (MMGBSA). We show for the first time the benefits of physics-informed pre-training to train a surrogate MMGBSA model on a collection of over 1.4 million 3D trajectories collected from MD simulations of the CASF-2016 benchmark. SurGBSA demonstrates a dramatic 6,497 speedup versus a traditional physics-based single-point MMGBSA calculation while nearly matching single-point MMGBSA accuracy on the challenging pose ranking problem for identification of the correct top pose ( difference). Our work advances the development of molecular foundation models by showing model improvements when training on MD simulations. Models, code and training data are made publicly available.
@@
In this Article, we present Surrogate mmGBSA (SurGBSA) for binding 3D structure ranking. Our SurGBSA method comprises (1) the first study of surrogate MMGBSA (or MMPBSA) models trained on snapshot level energetics with an order of magnitude higher snapshot resolution than previously studied and (2) demonstrate a significant improvement in learning the MMGBSA energetics due to physics-informed pretraining. We give the first practical evaluation of MMGBSA surrogate models based on docking pose ranking performance, compared to both the ensemble-averaged MMGBSA and single-point MMGBSA baseline methods. Additionally, we (3) investigate a range of molecular graph encoder modules (GNN, EGNN, EGMN) and their learned representations in detail. Our best-performing SurGBSA model achieves an approximate 6,497 speedup in pose re-ranking while preserving comparable ranking performance to the most efficient MMGBSA baseline, with the difference in correct pose identification being a single sample (). This dramatic efficiency gain enables the high-throughput generation of physics-informed pose ranking, providing the computational foundation necessary for training large-scale MD models that could accelerate virtual screening and drug discovery pipelines. Additionally, in contrast to ensemble-averaged MMGBSA based upon MD, SurGBSA can perform inference without the need to perform additional simulation, greatly reducing the computational resources required for virtual screening. SurGBSA can thus allow for rapid pose ranking comparable to single-point MMGBSA with a dramatic increase in computational efficiency while learning a robust molecular representation. Lastly, we make our implementation and data publicly accessible to support future work in MD foundation model development.
@@ 2.1 Learning a free-energy surrogate model from MD
In Figure 4 and Table 2 we detail the comparison in docking pose re-ranking with respect to the MMGBSA score based on all docking pose MD snapshots versus SurGBSA, using the various molecular encoder networks considered in this work, and the SP-GBSA baseline. Although our results show that SP-GBSA achieves the highest mean Spearman taken over all PDB structures in each test set, the difference in this ranking performance versus our best SurGBSA model is . Compared to using our baseline GNN encoder module, the improvement by our best SurGBSA model is significant at . Lastly, in Table 3 we give the results for correct pose identification. For each PDB structure, we presented each model with the five corresponding docking poses we generated. The poses are then ranked according to both the model prediction and to their ensemble-averaged MMGBSA energies. If the model top ranking agrees with the MMGBSA ranking, then we count this as a success. We again note a trend in terms of model complexity with our best overall method using the physics-informed EGMN extended training run. When comparing to the baseline re-scoing method, SP-GBSA, the difference is only a single test sample.
@@ 2.2 Efficiency of SurGBSA
| Method | Host | Scoring (s) | Speedup |
| MMGBSA [37] | CPU (Dane) | 11.079 | 0.2 |
| SP-GBSA [37] | CPU (Dane) | 1.949 | 1.0 |
| SurGBSA | AMD MI300A (Tuolumne) | 0.002 | 974.5 |
| SurGBSA | Nvidia H100 (Matrix) | 0.0003 | 6496.7 |
The MMGBSA score itself is relatively expensive compared to docking scoring functions as it requires MD simulation to sample conformations to improve its accuracy [26]. In Section 2.1 we discussed the ability to accurately recapitulate the MMGBSA score through a physics-informed surrogate model. Next, the surrogate model’s computational efficiency is compared to the MMGBSA score. In Table 4 we provide performance analysis collected on HPC clusters (Dane, Matrix, and Tuolumne). We considered two versions of MMPBSA.py [37] included in the Amber 22 MD simulation toolkit, a single core implementation and a multithreaded MPI implementation (MMPBSA.MPI.py). We only considered the Generalized Born (GB) solvation model as it is more efficient and serves as a more competitive baseline while scaling reasonably to the full dataset we consider in this work. As the MMGBSA calculation is parallelized over the frame dimension, SP-GBSA can only take advantage of a single core of the CPU per frame. Our results show that for the latency required to score a single docking pose versus the most competitive baseline, SP-GBSA, our SurGBSA model can achieve a speed-up improvement of approximately 6,497. Compared to running the parallel MPI version of the MMGBSA scoring function, using all snapshots from an MD trajectory, the speedup is approximately 36,930. Given the improved efficiency in conjunction with docking pose ranking performance as compared to the SP-GBSA and MMGBSA baselines, we believe that SurGBSA is capable of serving as a scalable method for docking pose re-scoring that can be used to supplement more expensive methods, including MMGBSA based on MD, by providing an efficient and comparatively accurate predictor. SurGBSA does not require MD for inference; thus, it can be applied to significantly larger datasets of docking poses by several orders of magnitude, improving the size of the tractable search space. This scalability allows SurGBSA to augment the more expensive MD-based scoring methods as a pre-processing step for ranking 3D structures of protein-ligand co-complexes.
@@ 3 Conclusion
SurGBSA is a new surrogate model for physics-based scoring of binding affinity of protein-ligand co-complex structures. Extending previous physics-informed pre-training, the new model learns an accurate and efficient mapping between molecular structure phase space and the corresponding binding free energy surface using the MMGBSA score. We validate our approach using a rigorous train/test split of representative protein-ligand structures and cross-validation with multiple random seeds for each evaluated model architecture. Results demonstrate that the SurGBSA model, which combines ProtMD[27] physics-informed pre-training with learning a mapping between molecular conformation and binding free energy produces a more robust representation compared to methods not trained on molecular dynamics trajectories. The SurGBSA model differs from the baseline physics-based single point (SP) GBSA calculation by a single ranking (i.e. ), when evaluated on a challenging pose ranking problem, with approximately a 6,497 speedup.
@@ 4.1 Dataset
To prepare the dataset for learning, we filter the simulation frames based upon the pocket-ligand co-complex structure to the starting frame, excluding all frames beyond to account for instabilities in the simulations while retaining meaningful fluctuations and conformational changes. We also filter out frames with MMGBSA scores greater than or equal to -10 and with MMGBSA scores less than -100. We then scaled the MMGBSA scores by dividing by a constant factor of -10, as we found that it leads to better numerical stability during training. We only keep the atoms within a radius of 6 angstroms from the ligand and use a maximum of 600 atoms while we keep the full ligand in each frame. We used the MDAnalysis [39, 40] python library to process the MD simulations. We aligned each trajectory with the respective crystal structure simulation and collected the RMSD for each frame with respect to the crystal pose. We filter out all water molecules, metals, and ions. To measure generalization performance, we considered a K-fold split of the core set complexes with , the default value of the scikit-learn utility, resulting in approximately 80% of the data used for training and 20% for testing. In total, after filtering based on the MMGBSA score and pocket-ligand co-complex starting point RMSD, we retain a total of 1,070,272 frames of simulation out of the 1.4 million collected from the MD simulations (). We encode the atom types and positions as input for the models.
@@ 4.2 MMGBSA
The MMGBSA method is an end-point free energy method that is more computationally efficient compared to Alchemical free energy (AFE) methods while being more accurate than docking scoring functions [26]. Due to these properties, MMGBSA has found application for evaluating docking poses and predicting binding affinities [26]. For our baselines, we use the MMPBSA.py script provided in the AmberTools22 software package [41]. We calculated single-point MMGBSA scores for each simulation frame in our dataset. We formed two distinct MMGBSA baselines; (1) SP-GBSA: uses the docked or crystal pose and the corresponding single-point MMGBSA score, and (2) MMGBSA which is based on the average of the single-point MMGBSA scores for each docking pose or crystal structure MD trajectory. SP-GBSA serves to function as a realistic efficient approximation to conserve computational resources when selecting ligands and their poses for further MD simulation. MMGBSA serves to represent the most accurate version of the re-scoring function based on all available conformational dynamics.
@@ 4.7 Model training and optimization
We considered a grid search over various settings of the hyperparameters. We considered batch sizes of 64, 128, and 256. We considered learning rates of and . We train all models using the AdamW optimizer with no weight decay in order to constrain the hyperparameter optimization and focus our study on the learning dynamics induced by the choice of learning rate and batch size instead. We considered 5 distinct random splits of the CASF-2016 complexes, which represent a representative sub-sampling of the PDBBind 2019 refined set structures. We trained a model, independently, for each unique configuration of hyperparameters on each of the 5 folds. Further, we replicated each experiment using 3 random seeds, thus for each configuration of hyperparameters we have a total of 15 independent training runs. In total, we conducted 5400 training runs for our study using the PyTorch [47] Distributed Data-Parallel (DDP) backend (v. 2.5.1), AMD ROCm 6.2, on the Tuloumne HPC cluster equipped with the AMD MI300A Accelerated Processing Unit (APU)[48]. In contrast to the GPU accelerator where the CPU and GPU access disjoint memory spaces, the APU combines on-chip memory resources that are shared across both the CPU and GPU, improving memory bandwidth between the processing units[48]. For training runs using only static information (i.e. crystal structure and docking poses) we used 1 node with 4 total APUs per job. For training runs using the MD conformations, we used 4 nodes per job with a total of 16 APUs per job. We trained all models for a maximum of 100 epochs, regardless of the dataset. Using the best model configuration after 100 epochs, the ProtMD pre-trained EGMN encoder network, we extended training to 600 epochs and report this combination as our best SurGBSA model. We treat the graphs from different frames independently, i.e. as separate graphs and train all layers of the constituent molecular graph encoder and MMGBSA prediction head using the MSE loss. We report the results using the models with the hyperparameters that achieve the best average spearman correlation, averaged by PDB structure, on the test set.

### v2

@@ Abstract
Self-supervised pretraining from static structures of drug-like compounds and proteins enable powerful learned feature representations. Learned features demonstrate state of the art performance on a range of predictive tasks including molecular properties, structure generation, and protein-ligand interactions. The majority of approaches are limited by their use of static structures and it remains an open question, how best to use atomistic molecular dynamics (MD) simulations to develop more generalized models to improve prediction accuracy for novel molecular structures. We present SURrogate mmGBSA (SurGBSA) as a new modeling approach for MD-based representation learning, which learns a surrogate function of the Molecular Mechanics Generalized Born Surface Area (MMGBSA). We show for the first time the benefits of physics-informed pre-training to train a surrogate MMGBSA model on a collection of over 1.4 million 3D trajectories collected from MD simulations of the CASF-2016 benchmark. SurGBSA demonstrates a dramatic 27,927 speedup versus a traditional physics-based single-point MMGBSA calculation while nearly matching single-point MMGBSA accuracy on the challenging pose ranking problem for identification of the correct top pose ( difference). Our work advances the development of molecular foundation models by showing model improvements when training on MD simulations. Models, code and training data are made publicly available.
@@
In this Article, we present Surrogate mmGBSA (SurGBSA) for binding 3D structure ranking. Our SurGBSA method comprises (1) the first study of surrogate MMGBSA (or MMPBSA) models trained on snapshot level energetics with an order of magnitude higher snapshot resolution than previously studied and (2) demonstrate a significant improvement in learning the MMGBSA energetics due to physics-informed pretraining. We give the first practical evaluation of MMGBSA surrogate models based on docking pose ranking performance, compared to both the ensemble-averaged MMGBSA and single-point MMGBSA baseline methods. Additionally, we (3) investigate a range of molecular graph encoder modules (GNN, EGNN, EGMN) and their learned representations in detail. Our best-performing SurGBSA model achieves an approximate 27,927 speedup in pose re-ranking while preserving comparable ranking performance to the most efficient MMGBSA baseline, with the difference in correct pose identification being a single sample (). This dramatic efficiency gain enables the high-throughput generation of physics-informed pose ranking, providing the computational foundation necessary for training large-scale MD models that could accelerate virtual screening and drug discovery pipelines. Additionally, in contrast to ensemble-averaged MMGBSA based upon MD, SurGBSA can perform inference without the need to perform additional simulation, greatly reducing the computational resources required for virtual screening. SurGBSA can thus allow for rapid pose ranking comparable to single-point MMGBSA with a dramatic increase in computational efficiency while learning a robust molecular representation. Lastly, we make our implementation and data publicly accessible to support future work in MD foundation model development.
@@ 2.1 Learning a free-energy surrogate model from MD
In Figure 4 and Table 2 we detail the comparison in docking pose re-ranking with respect to the MMGBSA score based on all docking pose MD snapshots versus SurGBSA, using the various molecular encoder networks considered in this work, and the SP-GBSA baseline. Although our results show that SP-GBSA achieves the highest mean Spearman taken over all PDB structures in each test set, the difference in this ranking performance versus our best SurGBSA model is . Compared to using our baseline GNN encoder module, the improvement by our best SurGBSA model is significant at . Lastly, in Table 3 we give the results for correct pose identification. For each PDB structure, we presented each model with the five corresponding docking poses we generated. The poses are then ranked according to both the model prediction and to their ensemble-averaged MMGBSA energies. If the model top ranking agrees with the MMGBSA ranking, then we count this as a success. We again note a trend in terms of model complexity with our best overall method using the physics-informed EGMN extended training run. Compared to the baseline scoring method SP-GBSA, the difference is only a single test sample.
@@ 2.2 Efficiency of SurGBSA
| Method | Host | Avg. Scoring (s) | Speedup |
| SP-GBSA [37] | CPU (Dane) | 8.378 | 1.0 |
| SurGBSA | AMD MI300A (Tuolumne) | 0.002 | 4,189.0 |
| SurGBSA | Nvidia H100 (Matrix) | 0.0003 | 27,926.7 |
The MMGBSA score itself is relatively expensive compared to docking scoring functions as it requires MD simulation to sample conformations to improve its accuracy [26]. In Section 2.1 we discussed the ability to accurately recapitulate the MMGBSA score through a physics-informed surrogate model. Next, the surrogate model’s computational efficiency is compared to the MMGBSA score. In Table 4 we provide performance analysis collected on HPC clusters (Dane, Matrix, and Tuolumne). We used the MMPBSA.py [37] implementation included in the Amber 22 MD simulation toolkit. We only considered the Generalized Born (GB) solvation model as it is more efficient and serves as a more competitive baseline while scaling reasonably to the full dataset we consider in this work. As the MMGBSA calculation is parallelized over the frame dimension, SP-GBSA can only take advantage of a single core of the CPU per frame. Our results show that for the latency required to score a single docking pose versus the most competitive baseline, SP-GBSA, our SurGBSA model can achieve a speed-up improvement of approximately 27,927. Given the improved efficiency in conjunction with docking pose ranking performance as compared to the SP-GBSA and MMGBSA baselines, we believe that SurGBSA is capable of serving as a scalable method for docking pose re-scoring that can be used to supplement more expensive methods, including MMGBSA based on MD, by providing an efficient and comparatively accurate predictor. SurGBSA does not require MD for inference; thus, it can be applied to significantly larger datasets of docking poses by several orders of magnitude, improving the size of the tractable search space. This scalability allows SurGBSA to augment the more expensive MD-based scoring methods as a pre-processing step for ranking 3D structures of protein-ligand co-complexes.
@@ 3 Conclusion
SurGBSA is a new surrogate model for physics-based scoring of binding affinity of protein-ligand co-complex structures. Extending previous physics-informed pre-training, the new model learns an accurate and efficient mapping between molecular structure phase space and the corresponding binding free energy surface using the MMGBSA score. We validate our approach using a rigorous train/test split of representative protein-ligand structures and cross-validation with multiple random seeds for each evaluated model architecture. Results demonstrate that the SurGBSA model, which combines ProtMD [27] physics-informed pre-training with learning a mapping between molecular conformation and binding free energy produces a more robust representation compared to methods not trained on molecular dynamics trajectories. The SurGBSA model differs from the baseline physics-based single point (SP) GBSA calculation by a single ranking (i.e. ), when evaluated on a challenging pose ranking problem, with approximately a 27,927 speedup.
@@ 4.1 Dataset
To prepare the dataset for learning, we filter the simulation frames based upon the pocket-ligand co-complex structure to the starting frame, excluding all frames beyond to account for instabilities in the simulations while retaining meaningful fluctuations and conformational changes. We also filter out frames with MMGBSA scores greater than or equal to -10 and with MMGBSA scores less than -100. We then scaled the MMGBSA scores by dividing by a constant factor of -10, as we found that it leads to better numerical stability during training. We only keep the atoms within a radius of 6 angstroms from the ligand and use a maximum of 600 atoms while we keep the full ligand in each frame. We used the MDAnalysis [39, 40] python library to process the MD simulations. We aligned each trajectory with the respective crystal structure simulation and collected the RMSD for each frame with respect to the crystal pose. We filter out all water molecules, metals, and ions. To measure generalization performance, we considered a K-fold split of the core set complexes with , the default value of the scikit-learn utility, resulting in approximately 80% of the data used for training and 20% for testing. In total, after filtering based on the MMGBSA score and pocket-ligand co-complex starting point RMSD, we retain a total of 1,070,272 frames of simulation out of the 1.4 million () collected from the MD simulations. We encode the atom types and positions as input for the models.
4.2 MD Simulations
We followed a standardized protocol for all MD simulations in this work, using the Amber22 [41] MD simulation software package. For each PDB structure, we used the conveyorLC [42] parallel docking program to generate ligand poses, selecting the top 5 poses for each PDB co-complex according to the Vina scoring function. Then for each pose, we used Sander to perform energy minimization for 1000 steps. The minimized structures were then heated to 300K under the NVT ensemble using a restraint on the protein backbone atoms of 1.0 kcal/mol for 200ps. The systems were then equilibrated to reach a pressure of 1 atm over 4.8 ns, using a weaker restraint on the atoms of 0.25 kcal/mol. Finally, we performed unrestrained production MD under the NPT ensemble for 10ns, with snapshots collected every 10 ps.
@@ 4.3 MMGBSA
The MMGBSA method is an end-point free energy method that is more computationally efficient compared to Alchemical free energy (AFE) methods while being more accurate than docking scoring functions [26]. Due to these properties, MMGBSA has found application for evaluating docking poses and predicting binding affinities [26]. For our baselines, we use the MMPBSA.py script provided in the AmberTools22 software package [41]. We calculated single-point MMGBSA scores for each simulation frame in our dataset. We formed two distinct MMGBSA baselines; (1) SP-GBSA: uses the docked or crystal pose and the corresponding single-point MMGBSA score, and (2) MMGBSA which is based on the average of the single-point MMGBSA scores for each docking pose or crystal structure MD trajectory. SP-GBSA serves to function as a realistic efficient approximation to conserve computational resources when selecting ligands and their poses for further MD simulation. MMGBSA serves to represent the most accurate version of the re-scoring function based on all available conformational dynamics. For the MMGBSA calculations, the Generalized Born implicit solvent model 5 was performed in pure water with salt contribution of 0 (igb=5, saltcon=0.0).
@@ 4.8 Model training and optimization
We considered a grid search over various settings of the hyperparameters. We considered batch sizes of 64, 128, and 256. We considered learning rates of and . We train all models using the AdamW optimizer with no weight decay in order to constrain the hyperparameter optimization and focus our study on the learning dynamics induced by the choice of learning rate and batch size instead. We considered 5 distinct random splits of the CASF-2016 complexes, which represent a representative sub-sampling of the PDBBind 2019 refined set structures. We treat the graphs from different frames independently, i.e. as separate graphs and train all layers of the constituent molecular graph encoder and MMGBSA prediction head using the MSE loss. We trained a model, independently, for each unique configuration of hyperparameters on each of the 5 folds. Further, we replicated each experiment using 3 random seeds, thus for each configuration of hyperparameters we have a total of 15 independent training runs. In total, we conducted 5400 training runs for our study using the PyTorch [48] Distributed Data-Parallel (DDP) backend (v. 2.5.1), AMD ROCm 6.2, on the Tuloumne HPC cluster equipped with the AMD MI300A Accelerated Processing Unit (APU) [49]. In contrast to the GPU accelerator where the CPU and GPU access disjoint memory spaces, the APU combines on-chip memory resources that are shared across both the CPU and GPU, improving memory bandwidth between the processing units [49]. For training runs using only static information (i.e. crystal structure and docking poses) we used 1 node with 4 total APUs per job. For training runs using the MD conformations, we used 4 nodes per job with a total of 16 APUs per job. We trained all models for a maximum of 100 epochs, regardless of the dataset. Using the best model configuration after 100 epochs, the ProtMD pre-trained EGMN encoder network, we extended training to 600 epochs and report this combination as our best SurGBSA model. Training our best model for 600 epochs took approximately 68.5 APU hours per device ( days). We report the results using the models with the hyperparameters that achieve the best average spearman correlation, averaged by PDB structure, on the test set.